# Memory-Efficient Sequential Pattern Mining with Hybrid Tries

### v2

@@ 2.2 Trie Models of the Data Set for SPM
For a sequential data set $\mathcal{SD}$, let $\mathcal{T}:=(\mathcal{N},\mathcal{A},\mathcal{L})$ be a labeled trie with node set $\mathcal{N}$, arc set $\mathcal{A}$, and a set of labels $\mathcal{L}$. The node set $\mathcal{N}$ is partitioned into $M+1$ subsets $\mathcal{N}_{0},\dots,\mathcal{N}_{M}$ referred to as layers. Layer $\mathcal{N}_{0}:=\{{\bf{r}}\}$ is a singleton containing auxiliary root node ${\bf{r}}$, and the remaining $j=1,\dots,M$ layers model the events of sequences $\mathcal{S}_{i}\in\mathcal{SD}$. Accordingly, each node $n\in\mathcal{N}\setminus\{{\bf{r}}\}$ is associated with an event label $e_{n}\in\mathcal{L}$ which denotes the event literal $e\in\mathbb{E}$ modeled by node $n$, and an itemset label $\mathbb{I}_{n}$ which denotes the position $j$ of itemset $\mathbb{I}_{i}^{j}$ in the sequence $\mathcal{S}_{i}$ modeled by path ${\mathbf{P}}$.
Using event and itemset labels $e_{n}$, a trie path ${\mathbf{P}}=({\bf{r}},n_{1},\dots,n_{|{\mathbf{P}}|})$ models the sequence $\mathcal{S}_{i}=\langle e_{n_{1}},\dots,e_{n_{|{\mathbf{P}}|}}\rangle$ belonging to itemsets $\langle\mathbb{I}_{n_{1}},\dots,\mathbb{I}_{n_{|{\mathbf{P}}|}}\rangle$. For notation convenience, we specify the sequence $\mathcal{S}_{i}$ modeled by a trie path using a transformation function $\mathbf{S}$, such that $\mathbf{S}({\mathbf{P}})=\mathcal{S}_{i}$.
The main advantage of trie models is their ability to model all overlapping prefixes $\mathcal{\check{S}}_{i}=\mathcal{\check{S}}_{j}$ of sequences in $\mathcal{SD}$ using only a single path ${\mathbf{P}}:\mathbf{S}({\mathbf{P}})=\mathcal{\check{S}}_{i}$. Accordingly, nodes $n\in\mathcal{N}\setminus\{{\bf{r}}\}$ are associated with a positive integer frequency label $f_{n}\in\mathcal{L}$ which denotes the number of prefixes $\mathcal{\check{S}}_{i}\in\mathcal{SD}$ modeled by path ${\mathbf{P}}=({\bf{r}},\dots,n)$. Example 4 illustrates the general trie model of the data set in our running example.
Figure 1: General trie model $\mathcal{T}$ of the data set in Table 1. Although the trie correctly models all sequences of the data set and their frequencies, its labels are insufficient for SPM.
@@
Forgoing the differences between web access mining and SPM, the trie models of Yang et al. (2007); Pei et al. (2000); Lu and Ezeife (2003); Rizvee et al. (2020) are identical in their node and arc set $(\mathcal{N},\mathcal{A})$. In fact, Lemma 5 shows that a trie model of a sequential data set is unique.
Lemma 5.
The trie model of a sequential data set is unique.
Proof.
Proof is by induction on the trie layers. The statement trivially holds for $\mathcal{N}_{0}=\{{\bf{r}}\}$. Now assume that the statement holds for a layer $\mathcal{N}_{j}:j<M$. By the definition of a trie, all nodes $n\in\mathcal{N}_{j+1}$ either have a unique parent $n\in\mathcal{N}_{j}$ or a unique event label $e_{n}$. Therefore, all nodes $n\in\mathcal{N}_{j+1}$ are unique. By the principle of mathematical induction, all layers $n_{j}\in\mathcal{T},j\in\{0,\dots,M\}$ are unique. Consequently, a trie model of a sequential data set is unique and of minimum size. ∎
The contribution of the trie models in the literature is the additional labels in $\mathcal{L}$ that are required to perform web access mining or SPM (see, for example, Pei et al. (2000); Lu and Ezeife (2003); Yang et al. (2007); Rizvee et al. (2020)). These label sets are critical in the mining process, and in the case of web access mining, do not generalize to SPM. This is due to the inherent difference between web access mining and SPM, where web access algorithms cannot distinguish between itemset and sequence extensions (Mabroukeh and Ezeife, 2010).
@@
TreeMiner traverses its trie model using the next-link matrices and mines patterns using the information stored in the parent-info and itemset labels. Although next-links enable faster traversal of the trie, and SPM is possible using the parent-info and itemset labels, storing them at each node of the trie model is costly in memory usage. In particular, Lemma 6 gives the worst-case space complexity of the TreeMiner algorithm.
@@
A node $n\in\mathcal{T}$ of the trie model of TreeMiner stores integer values $f_{n}$, $e_{n}$, $\mathbb{I}_{n}$, which are all bounded by $\mathcal{O}(\log(N))$, a next-link matrix bounded by $\mathcal{O}(|\mathbb{E}|^{2}(c^{\mathcal{V}}+\log(N)))$, a parent-info bitset bounded by $\mathcal{O}(c^{\mathcal{V}}+|\mathbb{E}|)$, and a children vector bounded by $\mathcal{O}\left(c^{\mathcal{V}}+|\mathbb{E}|\log(N)\right)$. The space complexity of a TreeMiner node is thus $\mathcal{O}\left(|\mathbb{E}|^{2}\log(N)\right)$.
By definition, a node $n$ of a trie model has at most $2|\mathbb{E}|$ children, an event $e\in\mathbb{E}$ representing an itemset extension and and an event $\bar{e}\in\mathbb{E}$ representing a sequence extension. The maximum number of children for a trie layer $\mathcal{N}_{j}$ is thus $\left(2|\mathbb{E}|\right)^{|\mathcal{N}_{j}|}$. As $|\mathcal{N}_{0}|=1$, the maximum number of nodes for any layer is $\left(2|\mathbb{E}|\right)^{M}$. On the other hand, we can have at most $N$ nodes at any layer of a trie model of a sequential data set $\mathcal{SD}$, where each sequence $\mathcal{S}_{i}\in\mathcal{SD}$ is modeled by exactly one path. The maximum number of nodes in a trie model is thus $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\right)$. For TreeMiner, this gives the overall worst-case memory complexity of its trie model as $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M|\mathbb{E}|^{2}\log(N)\right)$.
For its mining algorithm, TreeMiner stores one positional integer $j\leq M$ for at most every node in its trie, bounding its worst-case space complexity by $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\log(N)\right)$. The overall space complexity of TreeMiner is thus $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M|\mathbb{E}|^{2}\log(N)\right)$. Reasonably sized integers $N$ consume constant memory and reduce the complexity to
@@ 3 Binary and Hybrid Tries for Memory Efficient SPM
We introduce our binary trie $\mathcal{BT}$ in §3.1, and its corresponding mining algorithm in §3.2. We build on $\mathcal{BT}$ to develop a hybrid trie and its corresponding mining algorithm in §3.3.
@@ 3.1 Binary Tries
The main contribution of $\mathcal{BT}$ is its updated set of labels, which consumes a constant amount of memory. In addition to event, frequency, and itemset labels, $\mathcal{BT}$ stores at each node $n\in\mathcal{N}$ an additional ancestral label $a_{n}\in\mathcal{L}$, given by Definition 7
@@
As we later show, ancestor labels are sufficient to effectively mine all patterns from a trie model, and can be efficiently generated during its construction. Example 8 illustrates the binary-trie model for our running example.
@@
Consider the binary trie model of our running example given in Figure 3. The nodes $n_{2},n_{7},n_{8}$ are all siblings and model the children of their parent node $n_{1}$. Similarly, nodes $n_{3}$ and $n_{5}$ are siblings with parent node $n_{2}$.
@@ 3.2 The $\mathcal{BT}$Miner Algorithm
Given a data set $\mathcal{SD}$, the SPM task involves finding all patterns $\mathcal{P}$ such that $supp(\mathcal{P})\geq\theta\times N$. In prefix projection, the algorithm is initialized by frequent patterns containing a single event, that is, $\mathcal{P}=\langle\bar{e}\rangle:e\in\mathbb{E},supp(e)\geq\theta\times N$. At each subsequent iteration, a pattern $\mathcal{P}$ is extended by a single event $e\in\mathbb{E}$ to $\langle\mathcal{P},e\rangle$, and checked to satisfy $supp(\langle\mathcal{P},e\rangle)\geq\theta\times N$. A pattern extension can be categorized into itemset extension or sequence extension. In an itemset extension, a pattern $\mathcal{P}$ is extended to $\langle\mathcal{P},e\rangle$, extending its last itemset $\mathbb{I}^{|\mathcal{P}|}$ by a single event. In a sequence extension, a pattern is extended to $\langle\mathcal{P},\bar{e}\rangle$, extending $\mathcal{P}$ by a new itemset $\mathbb{I}=\{\bar{e}\}$.
@@
The proof follows from the minimal property of paths ${\mathbf{P}}=({\bf{r}},\dots,n):n\in\mathcal{N}(\mathcal{P})$. Assume by contradiction that the prefix $\mathcal{\check{S}}_{i}:\mathcal{P}\sqsubseteq\mathcal{\check{S}}_{i}$ modeled by a minimal path ${\mathbf{P}}:\mathbf{S}({\mathbf{P}})=\mathcal{\check{S}}_{i}$ is not of minimum size. Then there must exist another prefix $\mathcal{\check{S}}^{\prime}_{i}:\mathcal{\check{S}}^{\prime}_{i}\sqsubset\mathcal{\check{S}}_{i},\mathcal{P}\sqsubseteq\mathcal{\check{S}}^{\prime}_{i}$. Let ${\mathbf{P}}^{\prime}=({\bf{r}},\dots,n^{\prime})$ be the trie path that models prefix $\mathcal{\check{S}}^{\prime}_{i}$. As trie paths are unique and $\mathcal{\check{S}}^{\prime}_{i}\sqsubset\mathcal{\check{S}}_{i}$, we have ${\mathbf{P}}=\langle{\mathbf{P}}^{\prime},\dots,n$, a contradiction to the minimality of path ${\mathbf{P}}$. As node set $\mathcal{N}(\mathcal{P})$ contains all nodes $n:\mathcal{P}\sqsubseteq\mathbf{S}(({\bf{r}},\dots,n))$, then it models all prefixes $\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}:\mathcal{P}\sqsubseteq\mathcal{\check{S}}_{i}$ for all $\mathcal{S}_{i}\in\mathcal{SD}$. ∎
@@
For the converse, assume by contradiction that $\mathbb{I}_{n^{\prime}}\neq\mathbb{I}_{n}$ and $\mathbb{I}_{a_{n^{\prime}}}\leq\mathbb{I}_{n}$, but that the path ${\mathbf{P}}$ is not minimal. Then path ${\mathbf{P}}$ must be of the form ${\mathbf{P}}=(n,\dots,n^{\prime\prime},\dots,n^{\prime})$, where ${\mathbf{P}}^{\prime}=(n,\dots,n^{\prime\prime}):\langle\mathcal{P},\bar{e}_{n^{\prime}}\rangle\sqsubseteq\mathbf{S}\left((n,\dots,a_{n^{\prime\prime}})\right)$ is minimal. As $n^{\prime\prime}$ corresponds to a sequence extension $\langle\mathcal{P},\bar{e}_{n^{\prime}}\rangle$, we must have $e_{n^{\prime\prime}}=e_{n^{\prime}},\mathbb{I}_{n^{\prime\prime}}>\mathbb{I}_{n}$. This contradicts $\mathbb{I}_{a_{n^{\prime}}}\leq\mathbb{I}_{n}$, by definition of ancestor label $a_{n^{\prime}}$.
@@
    $\exists e_{j}\in\mathbb{I}^{|\mathcal{P}|}:\not\exists n^{\prime\prime}\in{\mathbf{P}}:e_{n^{\prime\prime}}=e_{j},\mathbb{I}_{n^{\prime\prime}}=\mathbb{I}_{n^{\prime}}$, or

### v3

@@ 2.2 Trie Models of the Data Set for SPM
For a sequential data set $\mathcal{SD}$, let $\mathcal{T}:=(\mathcal{N},\mathcal{A},\mathcal{L})$ be its labeled trie model with node set $\mathcal{N}$, arc set $\mathcal{A}$, and a set of labels $\mathcal{L}$. The node set $\mathcal{N}$ can be partitioned into $M+1$ subsets $\mathcal{N}_{0},\dots,\mathcal{N}_{M}$, referred to as layers. Layer $\mathcal{N}_{0}:=\{{\bf{r}}\}$ is a singleton containing auxiliary root node ${\bf{r}}$, and the remaining $j=1,\dots,M$ layers model the events of sequences $\mathcal{S}_{i}\in\mathcal{SD}$. Accordingly, each node $n\in\mathcal{N}\setminus\{{\bf{r}}\}$ is associated with an event label $e_{n}\in\mathcal{L}$, which denotes the event literal $e\in\mathbb{E}$ modeled by node $n$; and an itemset label $\mathbb{I}_{n}$, which denotes the position $j$ of itemset $\mathbb{I}_{i}^{j}$ in the sequence $\mathcal{S}_{i}$ modeled by path ${\mathbf{P}}$.
Using event and itemset labels $e_{n}$, a trie path ${\mathbf{P}}=({\bf{r}},n_{1},\dots,n_{|{\mathbf{P}}|})$ models the sequence $\mathcal{S}_{i}=\langle e_{n_{1}},\dots,e_{n_{|{\mathbf{P}}|}}\rangle$ belonging to itemsets $\langle\mathbb{I}_{n_{1}},\dots,\mathbb{I}_{n_{|{\mathbf{P}}|}}\rangle$. For notation convenience, we specify the sequence $\mathcal{S}_{i}$ modeled by a trie path ${\mathbf{P}}$ using a transformation function $\mathbf{S}({\mathbf{P}})=\mathcal{S}_{i}$.
The main advantage of trie models is their ability to model all overlapping prefixes $\mathcal{\check{S}}_{i}$ of sequences in $\mathcal{SD}$ using only a single path ${\mathbf{P}}:\mathbf{S}({\mathbf{P}})=\mathcal{\check{S}}_{i}$. Accordingly, nodes $n\in\mathcal{N}\setminus\{{\bf{r}}\}$ are associated with a positive integer frequency label $f_{n}\in\mathcal{L}$ that denotes the number of prefixes $\mathcal{\check{S}}_{i}$ of sequences $\mathcal{S}_{i}\in\mathcal{SD}$ modeled by path ${\mathbf{P}}=({\bf{r}},\dots,n)$. Example 4 illustrates the general trie model of the data set in our running example.
Figure 1: General trie model $\mathcal{T}$ of the data set in Table 1. Although the general trie structure correctly models all sequences of the data set and their frequencies, its labels are insufficient for SPM.
@@
Forgoing the differences between web access mining and SPM, the trie models of Yang et al. (2007); Pei et al. (2000); Lu and Ezeife (2003); Rizvee et al. (2020) are identical in their node and arc set $(\mathcal{N},\mathcal{A})$. In particular, the contribution of the trie models in the literature is the additional labels in $\mathcal{L}$ that are required to perform web access mining or SPM (see, for example, Pei et al. (2000); Lu and Ezeife (2003); Yang et al. (2007); Rizvee et al. (2020)). These label sets are critical in the mining process, and in the case of web access mining, do not generalize to SPM. This is due to the inherent difference between web access mining and SPM, where web access algorithms cannot distinguish between itemset and sequence extensions (Mabroukeh and Ezeife, 2010).
@@
TreeMiner traverses its trie model using the next-link matrices and mines patterns using the information stored in the parent-info and itemset labels. Although next-links can enable faster traversal of the trie and SPM is possible using the parent-info and itemset labels, storing them at each node of the trie model is costly in memory usage. In particular, Lemma 5 gives the worst-case space complexity of the TreeMiner algorithm.
@@
A node $n\in\mathcal{T}$ of the trie model of TreeMiner stores integer values $f_{n}$, $e_{n}$, $\mathbb{I}_{n}$, which are all bounded by $\mathcal{O}(\log(N))$, a next-link matrix bounded by $\mathcal{O}((|\mathbb{E}|+1)c^{\mathcal{V}}+|\mathbb{E}|^{2}\log(N)))$, a parent-info bitset bounded by $\mathcal{O}(c^{\mathcal{V}}+|\mathbb{E}|)$, and a children vector bounded by $\mathcal{O}\left(c^{\mathcal{V}}+|\mathbb{E}|\log(N)\right)$. The space complexity of a TreeMiner node is thus $\mathcal{O}\left(|\mathbb{E}|^{2}\log(N)\right)$.
By definition, a node $n$ of a trie model has at most $2|\mathbb{E}|$ children, an event $e\in\mathbb{E}$ representing an itemset extension and an event $\bar{e}\in\mathbb{E}$ representing a sequence extension. The maximum number of children for a trie layer $\mathcal{N}_{j}$ is thus $\left(2|\mathbb{E}|\right)^{|\mathcal{N}_{j}|}$. As $|\mathcal{N}_{0}|=1$, the maximum number of nodes for any layer is $\left(2|\mathbb{E}|\right)^{M}$. On the other hand, we can have at most $N$ nodes at any layer of a trie model of a sequential data set $\mathcal{SD}$, where each sequence $\mathcal{S}_{i}\in\mathcal{SD}$ is modeled by exactly one path. The maximum number of nodes in a trie model is thus $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\right)$. For TreeMiner, this gives the overall worst-case memory complexity of its trie model as $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M|\mathbb{E}|^{2}\log(N)\right)$.
For its mining algorithm, TreeMiner stores one positional integer $j\leq M$ for at most every node in its trie, bounding its worst-case space complexity by $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\log(N)\right)$. Putting the two together, gives the overall space complexity of TreeMiner as $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M|\mathbb{E}|^{2}\log(N)\right)$. Reasonably sized integers $N$ consume constant memory and reduce the complexity to
@@ 3 Binary and Hybrid Tries for Memory Efficient SPM
We introduce our binary trie $\mathcal{BT}$ in §3.1, and its corresponding mining algorithm in §3.2. We build on $\mathcal{BT}$ to develop a hybrid trie $\mathcal{HT}$ and its corresponding mining algorithm in §3.3.
@@ 3.1 Binary Tries
The main contribution of $\mathcal{BT}$ is its updated set of labels which consumes a constant amount of memory. In addition to event, frequency, and itemset labels of a general trie model, $\mathcal{BT}$ stores at each node $n\in\mathcal{N}$ an additional ancestral label $a_{n}\in\mathcal{L}$, given by Definition 6.
@@
Intuitively, the ancestor label of a node $n$ tracks the first occurrence of event $e_{n}$ prior to node $n$, on the path $({\bf{r}},\dots,n)\in\mathcal{BT}$. As we later show, ancestor labels are sufficient to effectively mine all patterns from a trie model, and can be efficiently generated during its construction. Example 7 illustrates the binary-trie model for our running example.
@@
Consider the binary trie model of our running example given in Figure 3. The nodes $n_{2},n_{7},n_{8}$ are all siblings and model the children of their parent node $n_{1}$. Similarly, nodes $n_{3}$ and $n_{5}$ are siblings with parent node $n_{2}$. The ancestor of node $n_{7}$ is $a_{n_{7}}={\bf{r}}$ as there is no node $n^{\prime}:e_{n^{\prime}}=e_{n_{7}}$ on the path $({\bf{r}},n_{1},n_{7})$. The ancestor of node $n_{14}$ is $a_{n_{14}}=n_{3}$, as node $n_{3}$ is the first prior node to $n_{14}$ on the path $({\bf{r}},n_{1},n_{2},n_{3},n_{4},n_{14})$ with the same item $e_{n_{14}}=e_{n_{3}}$.
@@ 3.2 The $\mathcal{BT}$Miner Algorithm
Given a data set $\mathcal{SD}$, the SPM task involves finding all patterns $\mathcal{P}$ such that $supp(\mathcal{P})\geq\theta\times N$. In prefix-projection, the algorithm is initialized by frequent patterns containing a single event, that is, $\mathcal{P}=\langle\bar{e}\rangle:e\in\mathbb{E},supp(e)\geq\theta\times N$. At each subsequent iteration, a pattern $\mathcal{P}$ is extended by a single event $e\in\mathbb{E}$ to $\langle\mathcal{P},e\rangle$, and checked to satisfy $supp(\langle\mathcal{P},e\rangle)\geq\theta\times N$.
@@
The proof follows from the minimal property of paths ${\mathbf{P}}=({\bf{r}},\dots,n):n\in\mathcal{N}(\mathcal{P})$. Assume by contradiction that the prefix $\mathcal{\check{S}}_{i}:\mathcal{P}\sqsubseteq\mathcal{\check{S}}_{i}$ modeled by a minimal path ${\mathbf{P}}:\mathbf{S}({\mathbf{P}})=\mathcal{\check{S}}_{i}$ is not of minimum size. Then there must exist another prefix $\mathcal{\check{S}}^{\prime}_{i}:\mathcal{\check{S}}^{\prime}_{i}\sqsubset\mathcal{\check{S}}_{i},\mathcal{P}\sqsubseteq\mathcal{\check{S}}^{\prime}_{i}$. Let ${\mathbf{P}}^{\prime}=({\bf{r}},\dots,n^{\prime})$ be the trie path that models prefix $\mathcal{\check{S}}^{\prime}_{i}$. As trie paths are unique and $\mathcal{\check{S}}^{\prime}_{i}\sqsubset\mathcal{\check{S}}_{i}$, we have ${\mathbf{P}}=\langle{\mathbf{P}}^{\prime},\dots,n\rangle$, a contradiction to the minimality of path ${\mathbf{P}}$. As node set $\mathcal{N}(\mathcal{P})$ contains all nodes $n:\mathcal{P}\sqsubseteq\mathbf{S}(({\bf{r}},\dots,n))$, then it models all prefixes $\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}:\mathcal{P}\sqsubseteq\mathcal{\check{S}}_{i}$ for all $\mathcal{S}_{i}\in\mathcal{SD}$. ∎
@@
Proof.
By Proposition 8, all paths $({\bf{r}},\dots,n):n\in\mathcal{N}(\mathcal{P})$ model all minimum-sized prefixes $\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}:\mathcal{P}\sqsubseteq\mathcal{\check{S}}_{i}$ for all $\mathcal{S}_{i}\in\mathcal{SD}$. As paths $(n,\dots,n^{\prime})$ are minimal by definition and have $e_{n^{\prime}}=e$, then paths $({\bf{r}},\dots,n^{\prime})$ are also minimal and satisfy $\langle\mathcal{P},e\rangle\sqsubseteq\mathbf{S}\left(({\bf{r}},\dots,n^{\prime})\right)$. Finding all such paths for all nodes $n\in\mathcal{N}(\mathcal{P})$ thus constructs $\mathcal{N}(\langle\mathcal{P},e\rangle)$. ∎
@@
For the converse, assume by contradiction that $\mathbb{I}_{n^{\prime}}\neq\mathbb{I}_{n}$ and $\mathbb{I}_{a_{n^{\prime}}}\leq\mathbb{I}_{n}$, but that path ${\mathbf{P}}$ is not minimal. Then path ${\mathbf{P}}$ must be of the form ${\mathbf{P}}=(n,\dots,n^{\prime\prime},\dots,n^{\prime})$, such that ${\mathbf{P}}^{\prime}=(n,\dots,n^{\prime\prime}):\langle\mathcal{P},\bar{e}_{n^{\prime}}\rangle\sqsubseteq\mathbf{S}\left((n,\dots,a_{n^{\prime\prime}})\right)$ is minimal. As $n^{\prime\prime}$ corresponds to a sequence extension $\langle\mathcal{P},\bar{e}_{n^{\prime}}\rangle$, we must have $e_{n^{\prime\prime}}=e_{n^{\prime}},\mathbb{I}_{n^{\prime\prime}}>\mathbb{I}_{n}$. This contradicts $\mathbb{I}_{a_{n^{\prime}}}\leq\mathbb{I}_{n}$, by definition of ancestor label $a_{n^{\prime}}$.
@@
    $\exists e_{j}\in\mathbb{I}^{|\mathcal{P}|}:\nexists n^{\prime\prime}\in{\mathbf{P}}:e_{n^{\prime\prime}}=e_{j},\mathbb{I}_{n^{\prime\prime}}=\mathbb{I}_{n^{\prime}}$, or